OpenStudy (anonymous):

integral_(0)^inf t arctan t /(1+t^4) dt = ?

OpenStudy (zarkon):

are you sure you typed the problem correctly...it is a slight modification away from being a much nicer problem.

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} t \times \arctan(t) / (1+t^4) dt\]

OpenStudy (anonymous):

actually it is the same one that "no-data" has posted, i just substitute t=tan (x/2), however, wolframalpha is able to calculate this !

OpenStudy (zarkon):

there is no need for the Weierstrass substitution to do that problem.